# The First and Second-Order Asymptotics of Covert Communication over AWGN Channels

### v5

@@ III-D Converse Bound
The above corollary show that there exists “part” of the whole codewords such that the maximal power is bounded. By using existing converse bounds under maximal power constraint, we can upper bound the size of that part, and hence the size of each subcode.
@@
For AWGN covert channels with channel models in (5) and (6), we have
$\log$ $M^{*}(n,\epsilon,\sigma,\delta)\leq(n\delta\log e)^{1/2}\sigma^{2}$
$+$ $(2)^{1/2}(n\delta)^{1/4}\sigma(\log e)^{3/4}\cdot Q^{-1}(\epsilon)+O(\log n).$ (141)
@@
- $\sqrt{n\mathsf{V}(\frac{n}{n-1}\Psi(\delta,\sigma,n))}\times\Phi^{-1}(1-\epsilon_{1})+O(\log n),$ (144)
@@
$\begin{split}\log&M^{*}(n,\epsilon,\sigma,\delta)\leq\frac{n}{2}\log\left(1+\frac{n}{n-1}\cdot\sqrt{\frac{4\eta\delta\ln 2}{n}}\cdot\sigma^{2}\right)\\ +&\sqrt{\frac{n\log^{2}e}{2}\cdot V_{2}(n)}\Phi^{-1}(\epsilon)+O(\log n),\end{split}$ (146)
where
$V_{2}(n)=\frac{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}}{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}+\frac{n}{4}}.$ (147)
@@
Note that both the first and second-order asymptotics of the converse bound in (7) differs from the achievability bound (6) in Theorem 6 on the coefficient $\mu$. It shows that the coding scheme in the achievability is optimal. In practice, the maximal throughput can be obtained by choosing proper parameter $\mu$ depending on the blocklength $n$ and the noise level $\sigma^{2}$ at the adversary.
@@
IV A Local Information Geometry Viewpoint of Gaussian Random Coding for Covert Communication
In Section III, truncated Gaussian distribution is adopted as the generating distribution of the AWGN covert codes. From the arguments in Section III-D, the distribution is optimal in the sense that it leads to nearly maximal coding rate when $n$ is large. In this section, we analyze it from a viewpoint of local information geometry, which helps us to understand the underlying principle of choosing generating distribution for random coding in covert communication.
@@ IV-A Quasi-$\varepsilon$-neighborhood Distributions
For any distribution $P_{1}(\bm{x})$ in $\epsilon$-neighborhood of $P_{0}(\bm{x})$ with $h(\bm{x})$ satisfying $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$, assume that $\int_{\mathbb{R}^{n}}p_{0}(\bm{x})h^{n}_{\varepsilon}(\bm{x})d\bm{x}<\infty$ for any $n\geq 2$, we have
@@ IV-B Realization of Quasi-$\varepsilon$-neighborhood Distributions
In the following, we reconsider the distribution family $P_{n}^{(n)}\sim\mathcal{N}(\bm{0},\mu\Psi(n)\bm{I}_{n})$ in Section II-E.
@@
From the above analysis, when the reference distribution is the distribution of noise, the truncated Gaussian distributions with decreasing variance are proper convolution operators that will lead to output distributions close to the distribution of noise. Hence, they are legitimate choice for the generating distributions of codebooks for AWGN covert channels. Note that the above conclusions holds for any $0<\tau<1$, while for covert communication over AWGN channels, the optimal $\tau$ is evaluated as $\frac{1}{2}$ in Theorem 6 and Theorem 7. The following section will explain the optimality of $\tau=\frac{1}{2}$ by investigating the asymptotic behavior of some divergences between Gaussian distributions.
V Square Root Law and Approximation of the Distribution of Noise
In Section III and III-D, we have witnessed that KL divergence between two Gaussian distributions with zero mean and i.i.d coordinates are the key ingredients in the evaluation of both the achievability and converse bounds of the throughput. In this section, the behavior of several divergences between two Gaussian distributions $\mathbb{P}^{(n)}_{1}\sim\mathcal{N}(\bm{0},(1+\theta_{n})\bm{I}_{n})$ and $\mathbb{P}^{(n)}_{0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ is investigated in Subsection A, where $\theta_{n}=c\cdot n^{-\tau}$ representing the decreasing power level. This provides more evidence of the connection between Square Root Law and the convergence behaviors of several divergences when $\tau\geq\frac{1}{2}$ and $\tau<\frac{1}{2}$, respectively. Our analysis provides an explanation why we choose $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $\tau=\frac{1}{2}$ in Gaussian random coding in the achievability.
V-A Asymptotic behaviors of $\mathbb{V}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ and $\mathbb{D}(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$
@@ V Square Root Law and Approximation of the Distribution of Noise
The following lemma is from the definition of Hellinger distance (3).
@@
If $\theta_{n}\sim c\cdot n^{-\tau}$ where $\sim$ represents the same asymptotic order, then
@@
If $\theta_{n}\sim c\cdot n^{-\tau}$ with $0<\tau<\frac{1}{2}$, then
@@
V-B Square Root Law and Approximation of the Distribution of Noise
@@ Appendix B Proof of Lemma 2
From the definition of $\bar{P}_{n}^{n}$ and $P_{n}^{n}$,
$\begin{split}\mathbb{D}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=&\int_{\bm{x}\in\Omega}\bar{g}^{(n)}(\bm{x})\log\frac{\bar{g}^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}d\bm{x}\\ =&\int_{\bm{x}\in\Omega}\bar{g}^{(n)}(\bm{x})\log\frac{1}{\Delta}d\bm{x}\\ =&-\log\Delta.\end{split}$ (163)
From the fact $\Omega=\mathfrak{B}^{n}_{0}(\sqrt{nP(n)})\backslash\mathfrak{B}^{n}_{0}(\sqrt{\mu^{2}nP(n)})$,
@@ Appendix E Proof of Lemma 10
$Q^{(n)}_{\bm{Z},\bm{C}_{1}}=\frac{1}{M}p_{1}+\frac{M-1}{M}p_{2},\,\ \ \ \,Q^{(n)}_{\bm{Z},\bm{C}_{2}}=\frac{1}{M}q_{1}+\frac{M-1}{M}q_{2}.$ (181)
@@ Appendix F Proof of Theorem 8
$\overset{(a)}{=}$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left(1+\varepsilon h_{\varepsilon}(\bm{x})\right)\left[\epsilon h_{\varepsilon}(\bm{x})-\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})\right]d\bm{x}$
$+$ $o(\varepsilon^{2})$
$=$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left[\varepsilon h_{\varepsilon}(\bm{x})+\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})-\frac{1}{2}\varepsilon^{3}h_{\varepsilon}^{3}(\bm{x})\right]d\bm{x}$
$+$ $o(\varepsilon^{2})$
@@
$\leq$ $\,\,\frac{1}{2}\varepsilon^{2}\log e+o(\varepsilon^{2})$
$\overset{(c)}{\leq}$ $\,\,\varepsilon^{2}\log e$ (185)
In the above inequality, we have used Taylor’s theorem in (a), Lemma 14 and the integrability assumption in (b) and $\varepsilon$ is a small number in (c).
@@ Appendix H Proof of Lemma 17
Let $\sigma_{1}^{2}=1+\theta_{n}$, we have [39]
$H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})=1-\left(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\right)^{\frac{n}{2}}.$
We just prove that $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ approaches 0 and $1$ with different $\tau$ as $n\rightarrow\infty$, respectively. The logarithm of $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ can be formulated as follows,
$\begin{split}&\frac{1}{2}n\ln\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\\ =&\frac{1}{2}n\ln\frac{2(1+\theta_{n})^{\frac{1}{2}}}{(2+\theta_{n})}\\ =&\frac{1}{2}n\left[\ln 2(1+\theta_{n})^{\frac{1}{2}}-\ln(2+\theta_{n})\right]\\ =&\frac{1}{4}n\ln\frac{4+4cn^{-\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}\\ \sim&-\frac{1}{4}\cdot\frac{c^{2}n^{1-2\tau}}{4cn^{-\tau}+4}.\end{split}$ (191)
  • When $\tau<\frac{1}{2}$, $1-2\cdot\tau>0$ and the above logarithm will approach -$\infty$ as $n\rightarrow\infty$. Consequently, $(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}$ approaches 0 as $n\rightarrow\infty$ and the conclusion is obtained.
  • When $\tau>\frac{1}{2}$, it is obvious that $\frac{1}{4}n^{1-2\tau}$ approaches 0 from the negative axis. Let $t=\left(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\right)^{\frac{n}{2}}$, from Taylor expansion, $e^{x}=1+x+o(x)$, we have
    $t\sim e^{-\frac{1}{4}n^{1-2\tau}}=1-\frac{1}{4}n^{1-2\tau}+o(\frac{1}{4}n^{1-2\tau})$
    and
    $\begin{split}&H(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})\\ =&\sqrt{1-t}\\ \sim&\sqrt{\frac{1}{4}n^{1-2\tau}+o(\frac{1}{4}n^{1-2\tau})}\\ \sim&\frac{1}{2}n^{\frac{1}{2}(1-2\tau)}.\end{split}$ (192)
    Thus, $H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ approaches 0 when $\tau>\frac{1}{2}$.

### v6

@@ III-D Converse Bound
The above corollary show that there exists “part” of the whole codewords such that the maximal power is bounded. By using existing converse bounds under maximal power constraint, we can upper bound the size of that part, and hence the size of each subcode, which lead to the converse part of Theorem 1.
@@
For covert communication over AWGN channels (5) and (6), any $(M,K,n,\epsilon,\sigma,\delta)$ code should satisfy (1).
@@
- $\sqrt{n\mathsf{V}(\frac{n}{n-1}\Psi(\delta,\sigma,n))}$
$\times$ $\Phi^{-1}(1-\epsilon_{1})+O(\log n),$ (143)
@@
$\begin{split}\log&M^{*}(n,\epsilon,\sigma,\delta)\\ \leq&\frac{n}{2}\log\left(1+\frac{n}{n-1}\cdot\sqrt{\frac{4\eta\delta\ln 2}{n}}\cdot\sigma^{2}\right)\\ +&\sqrt{\frac{n\log^{2}e}{2}\cdot V_{2}(n)}\Phi^{-1}(\epsilon)+O(\log n),\end{split}$ (145)
with
$V_{2}(n)=\frac{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}}{\sigma^{4}\delta\eta\ln 2+\sqrt{\sigma^{4}\delta\eta n\ln 2}+\frac{n}{4}},$ (146)
@@
IV On the Choice of Truncated Gaussian Distribution in Random Coding
In Section III-A, truncated Gaussian distribution is adopted as the generating distribution of the AWGN covert codes for achievability in Theorem 1. From the comparison between the achievability and the converse in Theorem 1, we can see that the generating distribution is optimal in the sense that it leads to nearly maximal coding rate when $n$ is large. In this section, the reason of choosing truncation Gaussian for random coding is investigated. We first explore it from a viewpoint of local information geometry, then we further analyze it by considering the asymptotic behavior of several divergences. The viewpoints in the section helps us to understand the underlying principle of choosing generating distribution for random coding in covert communication.
@@ IV-A Quasi-$\varepsilon$-neighborhood Distributions
For any sufficiently small $\varepsilon$, any distribution $P_{1}(\bm{x})$ in $\epsilon$-neighborhood of $P_{0}(\bm{x})$ with $h(\bm{x})$ satisfying $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$, assume that $\int_{\mathbb{R}^{n}}p_{0}(\bm{x})h^{n}_{\varepsilon}(\bm{x})d\bm{x}<\infty$ for any $n\geq 2$, we have
@@ IV-B Realization of Quasi-$\varepsilon$-neighborhood Distributions
In the following, we reconsider the distribution family $P_{n}^{(n)}\sim\mathcal{N}(\bm{0},\mu\Psi(n)\bm{I}_{n})$ in Section II-E and use $\|\cdot\|_{w}^{*}$ for the metric of weak convergence  [^3] Weak convergence with respect to bounded continuous functions. It has metric such as Lévy-Prokhorov metric[42].
@@
From the above analysis, when the reference distribution is the distribution of noise, the truncated Gaussian distributions with decreasing variance are proper convolution operators that will lead to output distributions close to the distribution of noise. Hence, they are legitimate choice for the generating distributions of codebooks for AWGN covert channels. The following section will explain the optimality of $\tau=\frac{1}{2}$ by investigating the asymptotic behavior of some divergences between Gaussian distributions.
IV-C Square Root Law and Approximation of the Distribution of Noise
In Section III, we have witnessed that KL divergence between two Gaussian distributions with zero mean and i.i.d coordinates are the key ingredients in the evaluation of both the achievability and converse bounds of the throughput. In this section, the behavior of several divergences between two Gaussian distributions $\mathbb{P}^{(n)}_{1}\sim\mathcal{N}(\bm{0},(1+\theta_{n})\bm{I}_{n})$ and $\mathbb{P}^{(n)}_{0}\sim\mathcal{N}(\bm{0},\bm{I}_{n})$ is investigated in Subsection A, where $\theta_{n}=c\cdot n^{-\tau}$ representing the decreasing power level. This provides more evidence of the connection between Square Root Law and the convergence behaviors of several divergences when $\tau\geq\frac{1}{2}$ and $\tau<\frac{1}{2}$, respectively. Our analysis provides an explanation why we choose $\Psi(n)=O(\frac{1}{n^{\tau}})$ with $\tau=\frac{1}{2}$ in Gaussian random coding in the achievability.
IV-C1 Asymptotic behaviors of $\mathbb{V}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})$ and $\mathbb{D}(\mathbb{P}^{(n)}_{1}\|\mathbb{P}^{(n)}_{0})$
@@ IV-C Square Root Law and Approximation of the Distribution of Noise
The following lemma is from the definition of the square of Hellinger distance (3).
@@
If $\theta_{n}\sim c\cdot n^{-\tau}$ with $0<\tau<1$ where $\sim$ represents the same asymptotic order, then
@@
If $\theta_{n}\sim c\cdot n^{-\tau}$ with $0<\tau<1$, then
@@
IV-C2 Square Root Law and Approximation of the Distribution of Noise
@@ Appendix B Proof of Lemma 2
Let $\Omega=\mathfrak{B}^{n}_{0}\big(\sqrt{n\Psi(n)}\big)\backslash\mathfrak{B}^{n}_{0}\big(\sqrt{\mu^{2}n\Psi(n)}\big)$, from the definition of $\bar{P}_{n}^{n}$ and $P_{n}^{n}$, we have $\frac{\bar{g}^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}=\frac{1}{\Delta}$ for $\bm{x}\in\Omega$, and $\bar{g}^{(n)}(\bm{x})=0$ for $\bm{x}\in\Omega^{c}$. Hence, we have
$\begin{split}\mathbb{D}(\bar{P}_{n}^{(n)},P_{n}^{(n)})=&\int_{\bm{x}\in\mathbb{R}^{n}}\bar{g}^{(n)}(\bm{x})\log\frac{\bar{g}^{(n)}(\bm{x})}{g^{(n)}(\bm{x})}d\bm{x}\\ =&\int_{\bm{x}\in\Omega}\bar{g}^{(n)}(\bm{x})\log\frac{1}{\Delta}d\bm{x}\\ =&-\log\Delta.\end{split}$ (162)
and
@@ Appendix E Proof of Lemma 10
$Q^{(n)}_{\bm{Z},\bm{C}_{1}}=\frac{1}{M}p_{1}+\frac{M-1}{M}p_{2},$
$Q^{(n)}_{\bm{Z},\bm{C}_{2}}=\frac{1}{M}q_{1}+\frac{M-1}{M}q_{2}.$ (180)
@@ Appendix F Proof of Theorem 9
$\overset{(a)}{=}$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left(1+\varepsilon h_{\varepsilon}(\bm{x})\right)$
$\times$ $\left[\epsilon h_{\varepsilon}(\bm{x})-\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})\right]d\bm{x}+o(\varepsilon^{2})$
$=$ $\log e\int_{\mathbb{R}^{n}}p_{0}(\bm{x})\left[\varepsilon h_{\varepsilon}(\bm{x})+\frac{1}{2}\varepsilon^{2}h_{\varepsilon}^{2}(\bm{x})\right$.
- $\left.\frac{1}{2}\varepsilon^{3}h_{\varepsilon}^{3}(\bm{x})\right]d\bm{x}+o(\varepsilon^{2})$
@@
$\overset{(c)}{\leq}$ $\,\,\frac{1}{2}\varepsilon^{2}\log e+o(\varepsilon^{2})$
$\overset{(d)}{\leq}$ $\,\,\varepsilon^{2}\log e$ (184)
In the above inequality, we have used Taylor’s theorem in (a), Lemma 14, $\underset{\bm{x}\in\mathbb{R}^{n}}{\sup}\,\,|h_{\varepsilon}(\bm{x})|\leq 1$ and the integrability assumption in in (b) and (c) and $\varepsilon$ is a small number in (d).
@@ Appendix H Proof of Lemma 17
Let $\sigma_{1}^{2}=1+\theta_{n}$, we have the square of Hellinger distance of two Gaussian distribution can be expressed as [39]
$H^{2}(\mathbb{P}^{(n)}_{1},\mathbb{P}^{(n)}_{0})=1-\left(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}\right)^{\frac{n}{2}}.$ (190)
To prove (158), we just need to prove
$(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})^{\frac{n}{2}}\rightarrow\begin{cases}1&\tau>\frac{1}{2},\\ 0&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty,$ (191)
which is equivalent to
$\frac{n}{2}\cdot\ln(\frac{2\sigma_{1}}{1+\sigma_{1}^{2}})\rightarrow\begin{cases}0&\tau>\frac{1}{2},\\ -\infty&0<\tau<\frac{1}{2},\end{cases}\,\ \,n\rightarrow\infty.$ (192)
The left-hand of (192) can be formulated as follows,
$\frac{1}{2}n\ln\frac{2\sigma_{1}}{1+\sigma_{1}^{2}}$
$=$ $\frac{1}{2}n\ln\frac{2(1+\theta_{n})^{\frac{1}{2}}}{(2+\theta_{n})}$
$=$ $\frac{1}{4}n\ln\bigg(\frac{2(1+\theta_{n})^{\frac{1}{2}}}{(2+\theta_{n})}\bigg)^{2}$
$=$ $\frac{1}{4}n\ln\frac{4(1+\theta_{n})}{(2+\theta_{n})^{2}}$
$=$ $\frac{1}{4}n\ln\frac{4+4cn^{-\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}$
$=$ $\frac{1}{4}n\ln\left[1-\frac{c^{2}n^{-2\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}\right]$
$\sim$ $-\frac{1}{4}\cdot\frac{c^{2}n^{1-2\tau}}{c^{2}n^{-2\tau}+4cn^{-\tau}+4}$ (193)
  • When $\tau>\frac{1}{2}$, it is obvious that the term (H) approaches 0 from the negative axis. Thus, we have proved the first part of (192).
  • When $\tau<\frac{1}{2}$, $1-2\tau>0$ and the term (H) approaches -$\infty$ as $n\rightarrow\infty$. Thus, the second part of (192) is proved.